OpenStudy (anonymous):

change trajectory of the path formula to make v0 the subject... please do this step by step

OpenStudy (experimentx):

please post the formula which is to be changed

OpenStudy (anonymous):

y=tan(theta)x -( (gx^2)/ (2(v0cos(theta))^2))

OpenStudy (experimentx):

y = ut sinQ - 1/2 g t x = ut cosQ i thought this was the formula, how did you get that?

OpenStudy (anonymous):

no thats uniform motion resnick book equation 4-25

OpenStudy (anonymous):

trajectory of the path...

OpenStudy (anonymous):

apparently its the horizontal motion and vertical motion put together....but i don't understand how they did that

OpenStudy (experimentx):

I understand what you are doing ... in my first equation, we have variable t, you get your equation by replacing t with x from second variable

OpenStudy (experimentx):

*** y = ut sinQ - 1/2 g t ^2

OpenStudy (experimentx):

and what do you mean by vo as subject .. is v. instantaneous velocity??

OpenStudy (anonymous):

yeh i kinda figured that a bit but then it still doesn't quite fit

OpenStudy (anonymous):

v0=u another name for initial velocity

OpenStudy (anonymous):

just curious, but does change trajectory of the path formula to make v0 the subject just mean "solve for v0"?

OpenStudy (anonymous):

yes

sam (.sam.):

\[y=\tan \theta x-(\frac{gx^2}{(2vcos(x))^2})\] Is it?

OpenStudy (anonymous):

so it is algebra you need right?

OpenStudy (anonymous):

yes .Sam but v is initial velocity

sam (.sam.):

\[v=\sqrt{\frac{gx^2}{4\cos^{2}x(\tan \theta x-y)}}\]

OpenStudy (anonymous):

yes and also how they put the equations together...the vertical motion and horizontal motion...

OpenStudy (anonymous):

\[y=x\tan (\theta) -\frac{gx^2}{2v_0cos^2(x)^2}\]

OpenStudy (anonymous):

or maybe i have it wrong and it is what .sam. wrote

OpenStudy (anonymous):

step by step would be good

OpenStudy (anonymous):

:)

OpenStudy (experimentx):

x and y is a particular point on your trajectory, theta is the angle of projection

OpenStudy (experimentx):

g is accn due to gravity, you have all parameters you need to calculate u

sam (.sam.):

There are some kinematic equations you can use, \[v_{f}=v_{i}+at\] \[v_{f}^2=v_{i}^2+2as\]

sam (.sam.):

horizontal has only one equation v=d/t

OpenStudy (anonymous):

how does those two equations make the trajectory of the path equation?

OpenStudy (anonymous):

y=tan(theta)x -( (gx^2)/ (2(v0cos(theta))^2)) \[y=(\tan \theta )x -gx^2/2(v0\cos \theta)^2 \]

sam (.sam.):

do you have the projectile question?

OpenStudy (anonymous):

no im just trying to understand the theory

OpenStudy (experimentx):

x, y givies you the path of the projectile ... thats parabolic

sam (.sam.):

because things mixed up pretty bad here y=tan(theta)x -( (gx^2)/ (2(v0cos(theta))^2))

OpenStudy (anonymous):

it says equation x-x0=(v0cos theta)t and y-y0=v0sin thetat-1/2gt^2 make the trajectory equation by replacing t with x....

OpenStudy (anonymous):

i did that but it doesn't quite change over

sam (.sam.):

So the idea is removing t from your new equation, x-x0=(v0cos theta)t .................(1) y-y0=v0sin thetat-1/2gt^2.......(2) From (1), \[ \huge t=\frac{x-x_{i}}{v_{i}\cos \theta}\] Sub (t) into (2), \[y-y_{i}=v_{i}\sin \theta t-\frac{1}{2}g t^2\] \[y-y_{i}=v_{i}\sin \theta (\frac{x-x_{i}}{v_{i}\cos \theta})-\frac{1}{2}g(\frac{x-x_{i}}{v_{i}\cos \theta})^2\]

sam (.sam.):

Did they say something like "solve in terms of" ?

OpenStudy (anonymous):

ahh i see it now....got mixed up with the change in x and y... well keep all the values there...i just want v0 to be the subject

OpenStudy (anonymous):

so i can sub all the values in sorry for slow reply, internet is playing up.

OpenStudy (anonymous):

because if its just y it is not that useful...the equation is so big

sam (.sam.):

\[\frac{1}{2}g(\frac{x-x_{i}}{v_{i}\cos \theta})^2=\tan \theta(x-x_{i})-(y-y_{i})\] \[g(\frac{x-x_{i}}{v_{i}\cos \theta})^2=2[\tan \theta(x-x_{i})-(y-y_{i})]\] \[\frac{(x-x_{i})^2}{v_{i}^2\cos^2 \theta}=\frac{2[\tan \theta(x-x_{i})-(y-y_{i})]}{g}\] \[\frac{1}{v_{i}^2}=\frac{2[\tan \theta(x-x_{i})-(y-y_{i})]\cos^2 \theta}{g(x-x_{i})^2}\] \[v_{i}^2=\frac{g(x-x_{i})^2}{2[\tan \theta(x-x_{i})-(y-y_{i})]\cos^2 \theta}\] \[v_{i}=\sqrt{\frac{g(x-x_{i})^2}{2[\tan \theta(x-x_{i})-(y-y_{i})]\cos^2 \theta}}\]

OpenStudy (anonymous):

ahhh i see what i did wrong....accidently forgot the square of cos^2 theta

OpenStudy (anonymous):

thanks